(2 pts) Calculate the equilibrium pressure of CO2 in the system at 3000C.
ΔH0 = 88.3 kJ and ΔS0 = 151.3 J/K. R = 8.314 J/mol K
PbCO3(s) ® PbO(s) + CO2(g)
ΔH0f(kJ/mol) -699.1 -217.3 -393.5
S0(J/mol K) 131 68.7 213.6
Answer #1

T = 3000 + 273 = 3273 K

ΔGo = ΔHo - T ΔSo

         = 88.3 - 3273 x 0.1513

         = -406.9

ΔGo   = - R T ln K

-406.9 = -8.314 x 10^-3 x 3273 x ln K

ln K = 14.95

K = 3.11 x 10^6

K= PCO2

PCO2   = 3.11 x 10^6 atm